OpenStudy (anonymous):

A bumper car with a mass of 200 kilograms traveling at a velocity of 3.5 meters/second collides with a stationary bumper car with a mass of 250 kilograms. If the two bumper cars stick and move together at a velocity of 2.3 meters/second, what is the kinetic energy of the system after the collision?

OpenStudy (anonymous):

just give me an equation

OpenStudy (anonymous):

summer school?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

humble tx

OpenStudy (anonymous):

hehe florida

OpenStudy (anonymous):

ok can you give me the equation or the answer

OpenStudy (anonymous):

KE = 1/2 mv^2 KE final = 1/2 (m1+m2) v^2 = 1/2 (200 + 250) (2.3)^2 = 1190 Joules